When a 17 gram(g) rubber stopper is placed in a beaker containing 25.46 mL of water, the water level rises to 38.00 mL. What is the density of the stopper in grams per mL?
Answer by Alan3354(69443)

When a 17 gram(g) rubber stopper is placed in a beaker containing 25.46 mL of water, the water level rises to 38.00 mL. What is the density of the stopper in grams per mL?
-----------------
38 - 25.46 = 12.54 ml change in water volume (not level).
The rubber stopper's volume is the change 12.54 ml
density = mass/vol = 17/12.54 =~ 1.356 gm/ml